When 2.22 g of a certain molecular compound X are dissolved in 75. g of benzene C6H6, the freezing point of the solution is measured to be 0.5 °C. Calculate the molar mass of X. If you need any additional information on benzene, use only what you find in the ALEKS Data resource. Also, be sure your answer has a unit symbol, and is rounded to the correct number of significant digits.
Answer #1

pure benzene freezing point = 5.50C

freezing point of solution = 0.50C

Tf = 5.5 - 0.5 = 5.0

Kf of benzene = 5.12 0C/m

we have a relation

Tf = Kf x m

m = molality = (W/MW) (1000 / mass of solvent in g)

Tf = Kf x W x 1000 / MW x mass of solvent

MW = Kf x W x 1000 / mass of solvent x Tf

MW = 5.12 x 2.22 x 1000 / 75 x 5.0

MW = 30.3 g/mol

answer = MW = 30.3 g/mol